OpenStudy (anonymous):

The function for the cost of materials to make a muffin is f(x) = Three fourths x + 3, where x is the number of muffins. The function for the selling price of those muffins is g(f(x)), where g(x) = 3x + 4. Find the selling price of 20 muffins.

OpenStudy (nurali):

Choice.

OpenStudy (anonymous):

f(x) = (3/4) x + 3 then x=20... first find f(x) then find g(f(x))....... then goon....

OpenStudy (anonymous):

18 49 51 58

OpenStudy (anonymous):

would it be 18

OpenStudy (nurali):

f(x) = (3/4) x + 3 g(x)=3x+4 selling price = g(f(x)) how to find/interpret g(f(x)) ?

OpenStudy (nurali):

reply

OpenStudy (anonymous):

No 58 is answer....

OpenStudy (anonymous):

can you explain it please

OpenStudy (anonymous):

First find f(x) = (3/4) x + 3 i.e., f(20) then insert that value in g(x)

OpenStudy (nurali):

g(f(x)) g(3/4x+3) 3(3/4x+3)+4 now plugin x=20

OpenStudy (anonymous):

given x=20 then f(20)=(3/4)(10)+3=18 then g(18)=3(18)+4=58 thats it.....

OpenStudy (anonymous):

ohh okay thank you @chetan552 and @Nurali if i plug it in 3*15=45+4=49

OpenStudy (anonymous):

x=20 then f(20)=(3/4)(20)+3=18

OpenStudy (nurali):

\[3(\frac{ 3 }{ 4 }(20)+3)+4\] \[3(3(5)+3)+4\] \[3(15+3)+4\] \[3(18)+4\] \[54+4\]

OpenStudy (nurali):

\[58\]

OpenStudy (anonymous):

no 3/4 *20 is 15

OpenStudy (anonymous):

many have done the math, 58 is your answer if you are looking for confirmation

OpenStudy (anonymous):

3/4 of 20 PLUS 3 = 18

OpenStudy (anonymous):

ohh okay thank you

OpenStudy (anonymous):

g(18) = 58

OpenStudy (nurali):

Anytime.

OpenStudy (nurali):

Welcome to Openstudy @anastasiad239